A rescue pilot drops a survival kit while her plane is flying at a height of 1,635 m with a horizontal velocity of 193 m/s. If air friction is ignored, how far in advance of the starving explorer’s drop zone should she release the package? (find horizontal displacement [range] of the package)
1
Expert's answer
2020-11-10T06:59:20-0500

Let's write the equations of motion of a survival kit (or the package) in horizontal and vertical directions:


"x=v_{0x}t, (1)""y=v_{0y}t+\dfrac{1}{2}gt^2, (2)"

here, "x" is the horizontal displacement of the package, "v_{0x}=193\ \dfrac{m}{s}" is the initial horizontal velocity of the package, "t" is the time that the package takes to reach the ground, "y=1635\ m" is the vertical displacement of the package, "v_{0y}=0\ \dfrac{m}{s}" is the initial vertical velocity of the package, "g=9.8\ \dfrac{m}{s^2}" is the acceleration due to gravity.

Let's first find the time that the package takes to reach the ground from the second equation:


"t=\sqrt{\dfrac{2y}{g}}."

Finally, we can substitute "t" into the first equation and calculate the horizontal displacement of the package:


"x=v_{0x}\sqrt{\dfrac{2y}{g}}=193\ \dfrac{m}{s}\cdot\sqrt{\dfrac{2\cdot 1635\ m}{9.8\ \dfrac{m}{s^2}}}=3525\ m."

Answer:

"x=3525\ m."